OpenStudy (anonymous):

One month bill rented 3 movies and 2 video games for a total of $17. The next month he rented 8 movies and 4 video games for a total of $37. Find the rental cost for each movie and each video game.

OpenStudy (anonymous):

3a+2a=17

OpenStudy (anonymous):

5.60 for 3 movies 8.50 for 2 video games 4.62 for 8 movies 9.25 for 4 video games each

OpenStudy (anonymous):

what exactly are u asking sweetie

OpenStudy (anonymous):

8a+4a=37

OpenStudy (anonymous):

that is incorrect @peachesncream she wants to know how much the bill for each move and video game cost

OpenStudy (anonymous):

i understand how u got the equation sillyseni but im not sure on how to figure out the answer

OpenStudy (radar):

System of equations. Use elimination process. m= movie cost v=video cost. 1st Eq: 3m + 2 v = $17.00 2nd Eq: 8m + 4v = $37.00 Now multiply 1st Eq by -2 getting -6m - 4v = -$34.00 Now add the 2nd eq to his getting 8m + 4v = $37 ---------------- 2m +0V =$3.00 solve for m as v has been eliminated m=$1.50 Now solve for v by substituting this value ($1.50) for m in the original equation.